# Finely Stratified Rerandomization Designs

### v2

@@ Abstract
We study estimation and inference on causal parameters under finely stratified rerandomization designs, which use baseline covariates to match units into groups (e.g. matched pairs), then rerandomize within-group treatment assignments until a balance criterion is satisfied. We show that finely stratified rerandomization does partially linear regression adjustment “by design,” providing nonparametric control over the stratified covariates and linear control over the rerandomized covariates. We introduce several new rerandomization designs, allowing for imbalance metrics based on nonlinear estimators. We also propose a novel minimax scheme that uses pilot data or prior information to minimize the computational cost of rerandomization, subject to a strict bound on statistical efficiency. While the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization is generically non-normal, we show how to restore asymptotic normality using ex-post linear adjustment tailored to the stratification. This enables simple asymptotically exact inference on superpopulation parameters, as well as efficient conservative inference on finite population parameters.
@@ 1 Introduction
Stratified randomization is commonly used to increase statistical precision in experimental research. [^2] For example, Cytrynbaum, (2023) reports a survey of 50 experimental papers in the AER and AEJ from 2018-2023, where 57% used some form of stratified randomization. Recent theoretical work (e.g. Bai et al., (2021)) has shown that fine stratification, which randomizes within small groups of units tightly matched on baseline covariate information, makes unadjusted estimators like difference of means semiparametrically efficient. [^3] See Cytrynbaum, (2024), Armstrong, (2022), and Bai et al., (2024) for more detailed discussion. In finite samples, however, the performance of such designs can deteriorate rapidly with the dimension of the stratification variables due to a curse of dimensionality in matching. [^4] Under regularity conditions, the convergence rate of finite sample variance to asymptotic variance is $O(n^{-2/(d+1)})$ for dimension $d$ covariates, see Cytrynbaum, (2024). This motivates the search for alternative designs that insist upon nonparametric balance for a few important covariates, but only attempt to balance linear functions of the remaining variables. In this paper, we study finely stratified rerandomization designs, which first tightly match the units into groups using a small set of important covariates, then rerandomize within-groups treatment assignments until a balance criterion on the remaining covariates is satisfied.
@@
Our first contribution is to derive the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization, allowing for estimation of generic causal parameters defined by moment equalities. We consider both superpopulation and finite population parameters, the latter of which may be more appropriate for experiments run in a convenience sample (Abadie et al., (2014)). As in previous work on rerandomization (e.g. Li et al., (2018)), the asymptotic distribution of GMM estimators is an independent sum of a normal and a truncated normal term. We show that, modulo this residual truncated term, the asymptotic variance of unadjusted estimation under stratified rerandomization is the same as that of semiparametrically adjusted GMM (e.g. Graham, (2011)) under an iid design. Intuitively, stratified rerandomization implements partially linear regression adjustment “by design.”
@@
Our second contribution is to introduce several novel forms of rerandomization based on nonlinear balance criteria. For example, we allow acceptance or rejection of an allocation based on the difference of covariate density estimates within each treatment arm, attempting to balance nonlinear features of the covariate distribution. Similarly, we propose a design that rerandomizes until a nonlinear estimate of the propensity score is approximately constant, effectively forcing the covariates to have no predictive power for treatment assignments. In both cases, these nonlinear rerandomization schemes are asymptotically equivalent to standard rerandomization based on a difference of covariate means, but with an implicit choice of covariates and acceptance region, which we characterize.
@@
Our third contribution is to study optimization of the balance criterion itself. We propose a novel minimax approach that allows the researcher to specify prior information about the relationship between covariates and outcomes, then rerandomizes until the worst case correlation between treatments and covariates consistent with this prior is small. We prove that this design minimizes the (asymptotic) computational cost of rerandomization, subject to a strict bound on statistical efficiency over the set of DGP’s consistent with the prior. If the prior information set contains the truth, this design strictly bounds the asymptotic variance within a small additive factor of the optimal semiparametrically adjusted variance. Extending this result, we show that if the prior information set is a confidence region estimated from pilot data, then this minimax design bounds the asymptotic variance in the main experiment with high probability.
@@
Our fourth contribution is to provide simple t-statistic and Wald based inference methods for general causal parameters under stratified rerandomization designs. To do this, we first characterize and provide a feasible implementation of the optimal ex-post linear adjustment for GMM estimation under stratified rerandomization. [^5] This extends recent work on optimal adjustment under pure stratified randomization for ATE estimation, e.g. see Cytrynbaum, (2023), Bai et al., (2023), or Liu and Yang, (2020). Crucially, optimal ex-post adjustment makes the asymptotic distribution insensitive to the rerandomization acceptance criterion, removing the truncated normal term from the limiting distribution and restoring asymptotic normality. For superpopulation parameters, our inference methods are asymptotically exact. For finite population parameters, our inference is conservative due to non-identification of the asymptotic variance, but still exploits the efficiency gains from both stratified rerandomization and ex-post optimal adjustment.
@@ 1.1 Related Literature
This paper builds on the literature on fine stratification in econometrics as well as the literature on rerandomization in statistics. Stratified randomization has a long history in statistics, see Cochran, (1977) for a survey. Recent work on fine stratification in econometrics includes Bai et al., (2021), Bai, (2022), Cytrynbaum, (2024), Armstrong, (2022), and Bai et al., (2024). Some important theoretical contributions to the literature on rerandomization include Morgan and Rubin, (2012) and Li et al., (2018), Wang et al., (2021), and Wang and Li, (2022). We build on both of these literatures, studying the consequence of rerandomizing treatments within data-adaptive fine strata. We show that finely stratified rerandomization does semiparametric (partially linear) regression adjustment “by design,” providing nonparametric control over a few important variables and linear control over the rest.
@@
For our main asymptotic theory (Section 3), the most closely related previous work is Wang et al., (2021) and Bai et al., (2024). Wang et al., (2021) study estimation of the sample average treatment effect (SATE) under stratified rerandomization, with quadratic imbalance metrics based on the Mahalanobis norm. We study rerandomization within data-adaptive fine strata, providing asymptotic theory for generic superpopulation and finite population causal parameters defined by moment equalities. We also allow for essentially arbitrary rerandomization acceptance criteria, not necessarily based on quadratic forms. Bai et al., (2024) study estimation of superpopulation parameters defined by moment equalities under pure stratified randomization. We extend these results to stratified rerandomization as well as generic finite population parameters, providing “SATE-like” versions of the parameters in Bai et al., (2024). [^6] These parameters can be seen as causal versions of the conditional estimand defined in Abadie et al., (2014). In concurrent work, Wang and Li, (2024) study GMM estimation of univariate superpopulation parameters under stratified rerandomization with fixed, discrete strata. We study significantly more general forms of stratification and rerandomization criteria than considered in their work, allowing for both finite and superpopulation parameters of arbitrary dimension and fine stratification with continuous covariates.
@@
For nonlinear rerandomization (Section 4), the closest related results are Ding and Zhao, (2024) and Li et al., (2021). Ding and Zhao, (2024) rerandomize based on the p-value of a logistic regression coefficient, while we rerandomize until a general smooth propensity estimate is close to constant. To the best of our knowledge, we present the first asymptotic theory for rerandomization based on the difference of nonlinear (e.g. density) estimates. For acceptance region optimization (Section 5), the closest related results are Schindl and Branson, (2024), who study the optimal choice of norm for quadratic rerandomization, while Liu et al., (2023) chooses a specific quadratic rerandomization using a Bayesian criterion, in both cases for rerandomization without stratification. We provide a novel minimax approach that accepts or rejects based on the value of a convex penalty function, tailored to prior information provided by the researcher. Our work on optimal adjustment (Section 6) extends recent work on adjustment for stratified designs, e.g. Liu and Yang, (2020), Cytrynbaum, (2023), Bai et al., (2023), to stratified rerandomization and GMM parameters. Finally our inference methods (Section 7) build on previous work by Abadie and Imbens, (2008), Bai et al., (2021), and Cytrynbaum, (2024). To the best of our knowledge we provide the first asymptotically exact inference for causal GMM parameters under stratified rerandomization, as well as conservative inference for their finite population analogues.
@@ 2 Framework and Designs
Consider data $W_{i}=(R_{i},S_{i}(1),S_{i}(0))$ with $(W_{i})_{i=1}^{n}\operatorname{\overset{iid}{\sim}}F$. The $S_{i}(d)\in\mathbb{R}^{d_{S}}$ denote potential outcome vectors for a binary treatment $d\in\{0,1\}$, while $R_{i}$ denote other pre-treatment variables, such as covariates. For treatment assignments $D_{i}\in\{0,1\}$, the realized outcome $S_{i}=S_{i}(D_{i})=D_{i}S_{i}(1)+(1-D_{i})S_{i}(0)$. In what follows, for any array $(a_{i})_{i=1}^{n}$ we denote $E_{n}[a_{i}]=n^{-1}\sum_{i=1}^{n}a_{i}$, with $\bar{a}_{1}=E_{n}[a_{i}D_{i}]/E_{n}[D_{i}]$ and $\bar{a}_{0}=E_{n}[a_{i}(1-D_{i})]/E_{n}[(1-D_{i})]$ Next, we define stratified rerandomization designs.
@@
    (Stratification). Partition the experimental units into $n/k$ disjoint groups $s$ with $\{1,\dots,n\}=\bigcup_{s}s$ disjointly and $|s|=k$. Let $\psi=\psi(R)$ with $\psi\in\mathbb{R}^{d_{\psi}}$ denote a vector of stratification variables, which may be continuous or discrete. Suppose the groups satisfy the homogeneity condition [^7] The matching condition in Equation 2.1 was introduced by Bai et al., (2021) for matched pairs randomization ($k=2$). See Bai, (2022) and Cytrynbaum, (2024) for generalizations.
@@
    (Check Balance). For rerandomization covariates $h=h(R)$, consider an imbalance metric $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})+o_{p}(1)$. [^8] In particular, we require $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})+o_{p}(1)$ under “pure” stratified randomization, the design in steps (1) and (2) only, studied e.g. in Cytrynbaum, (2024). We give several examples below. For an acceptance region $A\subseteq\mathbb{R}^{d_{h}}$, check if the balance criterion $\mathcal{I}_{n}\in A$ is satisfied. If so, accept $D_{1:n}$. If not, repeat from the beginning of (2).
Intuitively, steps (1) and (2) describe a data-adaptive “matched k-tuples” design, while step (3) rerandomizes within k-tuples until the balance criterion is satisfied. Equation 2.1 is a tight-matching condition, requiring that the groups are clustered locally in $\psi$ space. Cytrynbaum, (2024) provides algorithms to match units into groups that satisfy this condition for any fixed $k$.
Example 2.2 (Matched Pairs Rerandomization).
@@
Example 2.3 (Stratification).
Stratification without rerandomization can be obtained by setting $A=\mathbb{R}^{d_{h}}$ in Definition 2.1. Treatment effect estimation under such designs was studied in Bai, (2022), Cytrynbaum, (2024), and Bai et al., (2024). Definition 2.1 allows for fine stratification (also known as matched k-tuples), with the number of data-dependent groups $s=s(\psi_{1:n},\pi_{n})$ growing with $n$. It also allows for coarse stratification with fixed strata $T\in\{1,\dots,m\}$ and fixed $m$, as in Bugni et al., (2018), which can be obtained in this framework by setting $\psi=T$ and matching units into groups $s$ at random within each $\{i:T_{i}=k\}$.
Example 2.4 (Complete Randomization).
For $p=l/k$, we say that $D_{1:n}$ are completely randomized with probability $p$ if $P(D_{1:n}=d_{1:n})=1/\binom{n}{np}$ for all $d_{1:n}$ with $\sum_{i}d_{i}=np$ [^10] For notational simplicity, we may assume that $n=lk$ for some $l\in\mathbb{N}$. If so, we denote $D_{1:n}\sim\operatorname{CR}(p)$. Cytrynbaum, (2024) shows that $\operatorname{CR}(p)$ randomization can be obtained by setting $\psi=1$ and $A=\mathbb{R}^{d_{h}}$ in Definition 2.1, matching units into groups at random. Intuitively, random matched k-tuples is equivalent to complete randomization.
Causal Estimands. Next, we introduce a generic family of causal estimands defined by moment equalities. Let $g(D,R,S,\theta)\in\mathbb{R}^{d_{g}}$ be a score function for generalized method of moments (GMM) estimation. Recall $W=(R,S(1),S(0))$ and for $D|W\sim\operatorname{Bernoulli}(p)$ define $\phi(W,\theta)=E[g(D,R,S,\theta)|W]=pg(1,R,S(1),\theta)+(1-p)g(0,R,S(0),\theta)$. By construction, we have $E[\phi(W,\theta)]=0$ $\iff$ $E[g(D,R,S,\theta)]=0$. The function $\phi(W,\theta)$ provides a convenient parameterization to introduce our causal estimands.
@@
In what follows, we study GMM estimation of both $\theta_{0}$ and $\theta_{n}$ under stratified rerandomization designs, showing an asymptotic equivalence between stratified rerandomization and partially linear covariate adjustment. In particular, this framework allows us to introduce several useful finite population estimands $\theta_{n}$ that do not appear to have been considered previously in the literature. Note that GMM estimation of $\theta_{0}$ under pure stratification was studied in Bai et al., (2024) for the exactly identified case. Our finite population parameter $\theta_{n}$ can be viewed as a causal version of the finite population estimand defined in Abadie et al., (2014). [^11] See also the related finite population estimands studied under iid sampling and assignment in Xu, (2021) and Kakehi and Otsu, (2024).
Example 2.6 (ATE).
@@
For a more interesting example, consider the best parametric predictor of treatment effect heterogeneity in experiments with noncompliance.
@@
Let $D(z)$ be potential treatments for a binary instrument $z\in\{0,1\}$. Let $Y(d)$ be the potential outcomes, with realized outcome $Y=Y(D(Z))$. Suppose $D(1)\geq D(0)$, and define compliance indicator $C=\mathds{1}(D(1)>D(0))$, assuming $E[C]>0$. Imbens and Angrist, (1994) define the local average treatment effect $\operatorname{LATE}=E[Y(1)-Y(0)|C=1]$. Let $H=(Z-p)/(p-p^{2})$ and consider the score function $g(Z,D,Y,X,\theta)=(HY-HD\cdot f(X,\theta))\nabla_{\theta}f(X,\theta)$. Using standard LATE manipulations,
@@
Then $E[\phi(W,\theta)]=0$ is the first order condition of a treatment effect prediction problem in the complier population. In particular, for $\tau\equiv Y(1)-Y(0)$, the parameter $\theta_{0}$ is the best parametric predictor of treatment effects for compliers:
$\theta_{0}=\operatorname*{argmin}_{\theta}E[(\tau-f(X,\theta))^{2}|C=1].$
For example, if $Y$ is binary then $Y(1)-Y(0)\in\{-1,0,1\}$, so a scaled link function model $f(X,\theta)=2L(X^{\prime}\theta)-1$ may be appropriate. We can easily estimate marginal effects by adding $m(X_{i},\theta,\beta)=\beta-(\partial/\partial\theta^{\prime})f(X_{i},\theta)$ to the score function.
Example 2.8 (Finite Population Heterogeneity).
Continuing Example 2.7, note that for $\tau_{i}=Y_{i}(1)-Y_{i}(0)$ the corresponding finite population parameter is
@@
We can view $\theta_{n}$ as a “SATE-like” version of $\theta_{0}$, the best parametric predictor of treatment effects in the within-sample complier population. $\theta_{n}$ may be a more appropriate target for experiments run in a convenience sample. If $f(X,\theta)=X^{\prime}\theta$ linear, then $\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(\tau_{i}-X_{i}^{\prime}\theta)^{2}|C_{i}=1]$ is the within-sample best linear predictor. In the case of perfect compliance $C_{i}=1$ for all $i$, this is $\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(\tau_{i}-X_{i}^{\prime}\theta)^{2}]$, a finite-sample version of the best linear predictor of the conditional average treatment effect (CATE). The case $X=1$ recovers $\theta_{n}=E_{n}[Y_{i}(1)-Y_{i}(0)|C_{i}=1]$, the finite-population LATE, studied e.g. in Ren, (2023). Our inference methods in Section 7 produce tighter confidence intervals for these finite population parameters than $\theta_{0}$, since we only need to account for the uncertainty due to random assignment, with no sampling uncertainty.

### v3

@@ Abstract
We study estimation and inference on causal parameters under finely stratified rerandomization designs, which use baseline covariates to match units into groups (e.g. matched pairs), then rerandomize within-group treatment assignments until a balance criterion is satisfied. We show that finely stratified rerandomization does partially linear regression adjustment “by design,” providing nonparametric control over the stratified covariates and linear control over the rerandomized covariates. We introduce several new forms of rerandomization, allowing for imbalance metrics based on nonlinear estimators, and proposing a minimax scheme that minimizes the computational cost of rerandomization subject to a bound on estimation error. While the asymptotic distribution of GMM estimators under stratified rerandomization is generically non-normal, we show how to restore asymptotic normality using ex-post linear adjustment tailored to the stratification. We derive new variance bounds that enable conservative inference on finite population causal parameters, and provide asymptotically exact inference on their superpopulation counterparts.
@@ 1 Introduction
Stratified randomization is commonly used to increase statistical precision in experimental research. [^2] For example, Cytrynbaum, 2024a reports a survey of 50 experimental papers in the AER and AEJ from 2018-2023, where 57% used some form of stratified randomization. Recent theoretical work (e.g. Bai et al., (2021)) has shown that fine stratification, which randomizes treatment assignments within small groups of tightly matched units, makes unadjusted estimators like difference of means automatically semiparametrically efficient. [^3] See Cytrynbaum, 2024b , Armstrong, (2022), and Bai et al., 2024b for more detailed discussion. In finite samples, however, the performance of such designs can deteriorate rapidly with the dimension of the covariates used for stratification, due to a curse of dimensionality in matching. [^4] Under regularity conditions, the convergence rate of finite sample variance to asymptotic variance is $O(n^{-2/(d+1)})$ for dimension $d$ covariates, see Cytrynbaum, 2024b . This motivates the search for alternative designs that insist upon nonparametric balance for a few important covariates, but only attempt to balance linear functions of the remaining variables. In this paper, we study finely stratified rerandomization designs, which first tightly match the units into groups using a small set of important covariates, then rerandomize within-group treatment assignments until a balance criterion on the remaining covariates is satisfied.
@@
Our first contribution is to derive the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization, allowing for estimation of generic causal parameters defined by moment equalities. We consider both superpopulation and finite population parameters, the latter of which may be more appropriate for experiments run in a convenience sample (Abadie et al., (2014)), as is the case for the vast majority of experiments in economics (Niehaus and Muralidharan, (2017)). We introduce novel finite population parameters, studying a finite population local average treatment effect heterogeneity parameter in an application to Angrist et al., (2013). As in previous work on rerandomization (e.g. Li et al., (2018)), the asymptotic distribution of GMM estimators is an independent sum of a normal and a truncated normal term. We show that, modulo this truncated term, unadjusted GMM under stratified rerandomization behaves like semiparametrically adjusted GMM (e.g. Graham, (2011)) under an iid design, with automatic nonparametric control over the stratification covariates and linear control over the rerandomization covariates. Intuitively, stratified rerandomization implements partially linear regression adjustment “by design.”
@@
Our second contribution is to introduce novel forms of rerandomization based on nonlinear balance criteria. For example, we allow acceptance or rejection of an allocation based on the difference of estimated covariate densities between treatment and control units. We also study a design that rerandomizes until an estimated propensity score is approximately constant, forcing the covariates to have no predictive power for treatment assignments in our realized sample. We prove that the designs in a general family of nonlinear rerandomization methods are all asymptotically equivalent to standard rerandomization based on a difference of covariate means, with an implicit choice of covariates and balance criterion, which we characterize.
@@
Our third contribution is to study optimization of the balance criterion itself. We propose a novel minimax scheme that allows the researcher to specify prior information about the relationship between covariates and outcomes, then rerandomizes until the worst case covariate imbalance consistent with this prior is small. We prove that this design minimizes the (asymptotic) computational cost of rerandomization, subject to a strict bound on estimation error over the set of models consistent with the prior. If our prior information set contains the truth, then this design bounds the asymptotic variance of stratified rerandomization within a small additive factor of the optimal semiparametrically adjusted variance. If the information set is instead a Wald region estimated from pilot data, we show that our minimax design bounds the asymptotic variance in the main experiment with high probability.
@@
Our fourth contribution is to provide simple inference methods for generic causal parameters under stratified rerandomization designs. To do this, we first derive the optimal ex-post linear adjustment for GMM estimation, which depends on the stratification. [^5] This extends recent work on optimal adjustment under pure stratified randomization for ATE estimation, e.g. see Cytrynbaum, 2024a , Bai et al., 2024a , or Liu and Yang, (2020). Optimal adjustment makes the asymptotic distribution of GMM insensitive to the rerandomization acceptance criterion, removing the truncated normal term from the limiting distribution and restoring asymptotic normality. We also show that combining rerandomization with ex-post linear adjustment provides a form of double robustness to covariate imbalances, which helps explain the strong performance of this method in our simulations. For finite population causal parameters, the asymptotic variance is generically not identified (Neyman, (1990)). We derive novel identified variance bounds for general finite population causal parameters, enabling asymptotically conservative inference that still exploits the efficiency gains from both stratified rerandomization and optimal adjustment. For superpopulation parameters, we present new inference methods that are asymptotically exact.
Finally, we provide simulations and an empirical application to estimating treatment effect heterogeneity among compliers in Angrist et al., (2013), which both show the value of adding a rerandomization step to finely stratified designs. This effect can be seen clearly in Figure 3, which compares stratified rerandomization to stratification plus ex-post adjustment, and Figure 4, which compares stratified rerandomization to other designs like pure fine stratification. See the relevant sections below for more detailed discussion.
@@ 1.1 Related Literature
This paper builds on the literature on fine stratification in econometrics as well as the literature on rerandomization in statistics. Stratified randomization has a long history in statistics, see Cochran, (1977) for a survey. Recent work on fine stratification in econometrics includes Bai et al., (2021), Bai, (2022), Cytrynbaum, 2024b , Armstrong, (2022), and Bai et al., 2024b . A sample of some recent work in the statistics literature on rerandomization includes Morgan and Rubin, (2012) and Li et al., (2018), Wang et al., (2021), and Wang and Li, (2022). We build on both of these literatures, studying the consequence of rerandomizing treatments within data-adaptive fine strata. We show that finely stratified rerandomization does semiparametric (partially linear) regression adjustment “by design,” providing nonparametric control over a few important variables and linear control over the rest.
@@
For our main asymptotic theory (Section 3), the most closely related previous work is Wang et al., (2021) and Bai et al., 2024b . Wang et al., (2021) study estimation of the sample average treatment effect (SATE) under stratified rerandomization, with quadratic imbalance metrics based on the Mahalanobis norm. We study rerandomization within data-adaptive fine strata, providing asymptotic theory for generic superpopulation and finite population causal parameters defined by moment equalities. We also allow for essentially arbitrary rerandomization acceptance criteria, not necessarily based on quadratic forms. Bai et al., 2024b study estimation of superpopulation parameters defined by moment equalities under pure stratified randomization, without rerandomization. We extend these results to stratified rerandomization as well as generic finite population parameters, providing “SATE-like” versions of the parameters in Bai et al., 2024b . [^6] These parameters can be seen as causal versions of the conditional estimand defined in Abadie et al., (2014). In concurrent work, Wang and Li, 2024a study GMM estimation of univariate superpopulation parameters under stratified rerandomization with fixed, discrete strata. We study significantly more general forms of stratification and rerandomization criteria than considered in their work, allowing for both finite and superpopulation parameters of arbitrary dimension and fine stratification with continuous covariates.
@@
For nonlinear rerandomization (Section 4), the closest related results are Ding and Zhao, (2024) and Li et al., (2021). Ding and Zhao, (2024) rerandomize based on the p-value of a logistic regression coefficient, while we rerandomize until a general smooth propensity estimate is close to constant in $L_{2}$ norm. Li et al., (2021) simulate a density based rerandomization design, but provide limited theoretical results. To the best of our knowledge, we present the first asymptotic theory for rerandomization based on the difference of nonlinear (e.g. density) estimates. For acceptance region optimization (Section 5), the closest related results are Schindl and Branson, (2024), who study the optimal choice of norm for quadratic rerandomization, while Liu et al., (2023) chooses an optimal quadratic rerandomization design using a Bayesian criterion, in both cases for rerandomization without stratification. We propose a novel minimax approach that accepts or rejects based on the value of a convex penalty function, tailored to prior information provided by the researcher or estimated from a pilot.
Our work on optimal adjustment (Section 6) extends recent work on adjustment for stratified designs, e.g. Liu and Yang, (2020), Cytrynbaum, 2024a , Bai et al., 2024a , to stratified rerandomization and GMM parameters. Finally our work on inference under data-adaptive fine stratification (Section 7) builds on previous work by Abadie and Imbens, (2008), Bai et al., (2021), and Cytrynbaum, 2024b . Other recent work that has considered variance bounds for finite population causal parameters includes Aronow et al., (2014), Fogarty, (2018), Ding et al., (2019), Abadie et al., (2020), and Xu, (2021).
@@ 2 Framework and Designs
Consider data $W_{i}=(R_{i},S_{i}(1),S_{i}(0))$ with $(W_{i})_{i=1}^{n}\operatorname{\overset{iid}{\sim}}F$. The $S_{i}(d)\in\mathbb{R}^{d_{S}}$ denote potential outcome vectors for a binary treatment $d\in\{0,1\}$, while $R_{i}$ denote other pre-treatment variables, such as covariates. For treatment assignments $D_{i}\in\{0,1\}$, the realized outcome $S_{i}=S_{i}(D_{i})=D_{i}S_{i}(1)+(1-D_{i})S_{i}(0)$. In what follows, for any array $(a_{i})_{i=1}^{n}$ we denote $E_{n}[a_{i}]=n^{-1}\sum_{i=1}^{n}a_{i}$, with $\bar{a}_{1}=E_{n}[a_{i}|D_{i}=1]=E_{n}[a_{i}D_{i}]/E_{n}[D_{i}]$ and similarly $\bar{a}_{0}=E_{n}[a_{i}|D_{i}=0]$. Next, we define stratified rerandomization designs.
@@
    (Stratification). Partition the experimental units into $n/k$ disjoint groups (strata) $s$ with $\{1,\dots,n\}=\bigcup_{s}s$ disjointly and $|s|=k$. Let $\psi=\psi(R)$ with $\psi\in\mathbb{R}^{d_{\psi}}$ denote a vector of stratification variables, which may be continuous or discrete. Suppose the groups satisfy the matching condition [^7] The matching condition in Equation 2.1 was introduced by Bai et al., (2021) for matched pairs randomization ($k=2$). See Bai, (2022) and Cytrynbaum, 2024b for generalizations.
@@
    (Check Balance). For rerandomization covariates $h=h(R)$, consider an imbalance metric $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})+o_{p}(1)$. [^8] In particular, we require that $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})+o_{p}(1)$ under the law induced by “pure” stratified randomization, the design in steps (1) and (2) only, studied e.g. in Cytrynbaum, 2024b . For an acceptance region $A\subseteq\mathbb{R}^{d_{h}}$, check if the balance criterion $\mathcal{I}_{n}\in A$ is satisfied. If so, accept $D_{1:n}$. If not, repeat from the beginning of (2).
Intuitively, steps (1) and (2) describe a data-driven “matched k-tuples” design, while step (3) rerandomizes within k-tuples until the balance criterion is satisfied. Equation 2.1 is a tight-matching condition, requiring that the groups are clustered locally in $\psi$ space. Cytrynbaum, 2024b provides algorithms to match units into groups that satisfy this condition for any fixed $k$.
Example 2.2 (Pure Stratification).
Stratification without rerandomization can be obtained by setting $A=\mathbb{R}^{d_{h}}$ in Definition 2.1. Treatment effect estimation under such designs was studied in Bai, (2022), Cytrynbaum, 2024b , and Bai et al., 2024b . Definition 2.1 allows for fine stratification (also known as matched k-tuples), with the number of data-dependent groups $s=s(\psi_{1:n},\pi_{n})$ growing with $n$. It also allows for coarse stratification with strata $x\in\{1,\dots,m\}$ and fixed $m$, studied e.g. in Bugni et al., (2018). This can be obtained in our framework by setting $\psi=x$ and matching units into groups $s$ at random within each $\{i:x_{i}=k\}$.
Example 2.3 (Complete Randomization).
For $p=l/k$, we say that $D_{1:n}$ are completely randomized with probability $p$ if $P(D_{1:n}=d_{1:n})=1/\binom{n}{np}$ for all $d_{1:n}$ with $\sum_{i}d_{i}=np$ [^9] For notational simplicity, we may assume that $n=lk$ for some $l\in\mathbb{N}$. Equivalently, complete randomization is coarse stratification with $m=1$ above. This can be obtained by setting $\psi=1$ and $A=\mathbb{R}^{d_{h}}$ in Definition 2.1, matching units into groups at random.
Next, we discuss a convenient rerandomization scheme that allows the researcher to select the approximate number of draws until acceptance.
Example 2.4 (Mahalanobis Rerandomization).
Consider matched k-tuples rerandomization as in Equation 2.1. Define within-tuple demeaned covariates $\check{X}_{i}=X_{i}-k^{-1}\sum_{j\in s(i)}X_{j}$ and set $\Sigma_{n}=\operatorname{Var}(D)^{-1}\frac{k}{k-1}E_{n}[\check{X}_{i}\check{X}_{i}^{\prime}]$. Consider rerandomizing until
$n(\bar{X}_{1}-\bar{X}_{0})^{\prime}\Sigma_{n}^{-1}(\bar{X}_{1}-\bar{X}_{0})\leq\epsilon^{2}$ (2.2)
@@
Mahalanobis rerandomization uses a convenient choice of acceptance criterion, but the variance normalization and implicit acceptance region in Equation 2.4 are not generally efficient for estimating causal parameters. We provide alternative designs that optimize the shape of acceptance region $A$ in Section 5 below.
Causal Estimands. Next, we introduce a generic family of causal estimands defined by moment equalities. Let $g(D,R,S,\theta)\in\mathbb{R}^{d_{g}}$ be a score function for generalized method of moments (GMM) estimation. Recall $W=(R,S(1),S(0))$ and for $D|W\sim\operatorname{Bernoulli}(p)$ define $\phi(W,\theta)=E[g(D,R,S,\theta)|W]=pg(1,R,S(1),\theta)+(1-p)g(0,R,S(0),\theta)$. By construction, we have $E[\phi(W,\theta)]=0$ $\iff$ $E[g(D,R,S,\theta)]=0$. The function $\phi(W,\theta)$ provides a convenient parameterization to define our paired finite population and superpopulation causal estimands.
@@
In what follows, we study GMM estimation of both $\theta_{0}$ and $\theta_{n}$ under stratified rerandomization designs, showing an asymptotic equivalence between stratified rerandomization and partially linear covariate adjustment. In particular, this framework allows us to introduce several useful finite population estimands $\theta_{n}$ that do not appear to have been previously considered in the literature, such as Example 2.8 below. The estimand $\theta_{n}$ may be a more appropriate target for experiments run in a convenience sample, as is the case for the vast majority of experiments reported in the economics literature (Niehaus and Muralidharan, (2017)). Inference on $\theta_{n}$, provided in Section 7, is generically more powerful than for $\theta_{0}$, since we only have to account for estimation uncertainty due to random assignment, without extra variability from sampling into the experiment.
Remark 2.6 (Finite Population).
The parameter $\theta_{n}$ can be viewed as a causal version of the finite population estimand in Abadie et al., (2014), which they define in a regression setting with iid data. [^11] Also see the related finite population estimands in Xu, (2021) and Kakehi and Otsu, (2024). Their work only conditions on covariates $R$, so the asymptotic variance for estimating their finite population parameter is identified. By contrast, since we condition on $W=(R,S(1),S(0))$, the asymptotic variance for estimating $\theta_{n}$ is not identified, motivating the development of new variance bounds in Section 7 below.
Note also that GMM estimation of the superpopulation parameter $\theta_{0}$ under pure stratification was studied in Bai et al., 2024b .
Example 2.7 (ATE and SATE).
@@
Next, consider a setting where the researcher wants to estimate a parametric model of treatment effect heterogeneity in an experiment with noncompliance and randomized binary instrument $Z$. In the next example, we define finite population and superpopulation local average treatment effect (LATE) heterogeneity parameters, applying them in our empirical application to Angrist et al., (2013) below.
@@
Let $D(z)$ be potential treatments for a binary instrument $z\in\{0,1\}$. Let $Y(d)$ be the potential outcomes, with realized outcome $Y=Y(D(Z))$. Suppose $D(1)\geq D(0)$, and define compliance indicator $C=\mathds{1}(D(1)>D(0))$, assuming $E[C]>0$. Imbens and Angrist, (1994) define the $\operatorname{LATE}=E[Y(1)-Y(0)|C=1]$. Let $H=(Z-p)/(p-p^{2})$ and consider the score function $g(Z,D,Y,X,\theta)=(HY-HD\cdot f(X,\theta))\nabla_{\theta}f(X,\theta)$. Using standard LATE manipulations,
@@
The moment condition $E[\phi(W,\theta)]=0$ is the FOC of a treatment effect prediction problem in the complier population $C=1$. In particular, for $\tau\equiv Y(1)-Y(0)$, the parameter $\theta_{0}$ is the best parametric predictor $\theta_{0}=\operatorname*{argmin}_{\theta}E[(\tau-f(X,\theta))^{2}|C=1]$ of treatment effects for compliers. [^12] For example, if $Y$ is binary then $Y(1)-Y(0)\in\{-1,0,1\}$, so the link function model $f(X,\theta)=2L(X^{\prime}\theta)-1$ for $L=\text{Logit}$ may be appropriate. Specializing to $f(X,\theta)=X^{\prime}\theta$, this is the best linear predictor (BLP) of treatment effect heterogeneity among the compliers $\theta_{0}=\operatorname*{argmin}_{\theta}E[(\tau-X^{\prime}\theta)^{2}|C=1]$, while $f(X,\theta)=\theta$ recovers the $\operatorname{LATE}=E[\tau|C=1]$. Setting $E_{n}[\phi(W_{i},\theta)]=0$, the corresponding finite population parameter is
@@
We can also specialize to $f(X,\theta)=X^{\prime}\theta$ for a finite population version of the BLP of LATE. The finite population LATE was studied in Ren, (2023) under complete randomization, but the more general heterogeneity parameters here appear to be novel. We consider both $\theta_{0}$ and $\theta_{n}$ when studying treatment effect heterogeneity among compliers in the empirical application to Angrist et al., (2013) in Section 9.